OpenStudy (anonymous):

find the common ratio of the sequence -164,-82,-41,-20.5

OpenStudy (anonymous):

I need help ASAP please

OpenStudy (freckles):

divide term by previous term

OpenStudy (freckles):

if this is a geometric sequence it won't matter what term and previous term to that term you choose that is all of the following should be equal if this is indeed a geometric sequence: \[\frac{-82}{-164}=? \\ \frac{-41}{-82}=? \\ \frac{-20.5}{-41}=?\]